OpenStudy (anonymous):

Can anyone simplify: x^(1/3) x^(3/5)

OpenStudy (anonymous):

2x^(1/5)

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

it's:\[x^{\frac{1}{3}}x^{\frac{3}{5}}=x^{\frac{1}{3}+\frac{3}{5}}=x^{\frac{14}{15}}\]

hero (hero):

\[x^{1/3}x^{3/5} = x^{5/15}x^{9/15} = x^{14/15}\]

OpenStudy (anonymous):

that's what i thought it seemed too easy though. thanks!